OpenStudy (bluesea24):

what is difference btw sin5teta and 5sinteta

OpenStudy (harsimran_hs4):

1st is sin 5(tetha) i.e sin of five times theta 2nd is 5 sin(theta) i.e five times of sin theta

OpenStudy (bluesea24):

if value of tetha is 45 degree what is the ans in both cases

OpenStudy (harsimran_hs4):

1st one would be sin 5(45) = sin(135) = -1/sqrt(2) 2nd one is 5 sin(45) = 5 * (1/sqrt(2))

OpenStudy (bluesea24):

how sin(135)

OpenStudy (harsimran_hs4):

45 * 5 = 225 instead of 135 my bad you can now do rest of the calculations with 225 can you?

OpenStudy (bluesea24):

yes thanks pal......

OpenStudy (harsimran_hs4):

cool :)